# Sigma: Differential Rescaling of Query, Key and Value for Efficient Language Models

### v1

@@ 2.1 Formulation of DiffQKV Attention
Comparison with Existing Attention Mechanisms.
The three widely used attention mechanisms, MHA, Multi-Query Attention (MQA), and Grouped-Query Attention (GQA), can be viewed as special forms of the above algorithm. As shown in Figure 1, these three methods all employ the same dimensions for Q, K, and V heads ($d_{q}^{h}=d_{k}^{h}=d_{v}^{h}$). In terms of the head number, MHA sets the same number of Q, K, and V heads ($n_{q}^{h}=n_{k}^{h}=n_{v}^{h}$), MQA only employs a single K head and V head ($n_{k}^{h}=n_{v}^{h}=1$), and GQA features an intermediate number of K and V heads, yet still aligns their values ($n_{k}^{h}=n_{v}^{h}$). Additionally, most existing approaches directly retrieve the entire K and V caches before computing the attention scores, without tailored designed for LoadCacheK and LoadCacheV, respectively.
Research Questions.
Our goal is to find the optimal configurations for DiffQKV attention that facilitate strong model performance while ensuring inference efficiency by reducing memory consumption and burden on memory bandwidth. More specifically, the crucial configurations for DiffQKV attention include: (1) the number of QKV heads ($n_{q}^{h},n_{k}^{h},n_{v}^{h}$), (2) the dimensions of QKV heads ($d_{q}^{h},d_{k}^{h},d_{v}^{h}$), and (3) the strategies for loading K and V caches (LoadCacheK, LoadCacheV). The memory cost of KV cache per token is $O(n_{k}^{h}d_{k}^{h}+n_{v}^{h}d_{v}^{h})$. The time required to retrieve KV cache is determined by the practical data transfer conducted in LoadCacheK and LoadCacheV.
In the following section (§ 2.2), we present our experimental findings on how to set the above configurations, offering insights into achieving an optimal balance between model performance and inference efficiency. These insights lay the foundation for determining the architecture of Sigma.
2.2 Observations
In this section, we present our experimental findings on how various configurations of the Q, K, and V components in the DiffQKV attention would influence the model performance. To achieve this, we explore different model architectures and train them from scratch to assess their performance, entailing a substantial number of pre-training experiments. In the following experiments, we adopt 100 billion tokens from FineWeb-Edu (Penedo et al., 2024) as the pre-training data. All model architectures are scaled to approximately 1 billion parameters, which consists of 22-layer transformer blocks with a hidden dimension of 2048. To evaluate the model’s performance on different experimental settings, we employed the following benchmarks: HellaSwag (Hella.) (Zellers et al., 2019), OpenBookQA (ObQA) (Mihaylov et al., 2018), WinoGrande (Wino.) (Sakaguchi et al., 2021), ARC Challenge (ARC.) (Clark et al., 2018), PIQA (Bisk et al., 2020), SciQ (Welbl et al., 2017), BoolQ (Bool.) (Clark et al., 2019), LogiQA (Logi.) (Liu et al., 2020), and LAMBADA (Paperno et al., 2016)(LAMB.).
Observation 1
The model performance is more sensitive to the decrease in the number of V heads ($n^{h}_{v}$) than in the number of K heads ($n^{h}_{k}$). Specifically, we reduce the number of heads of K and V in the DiffQKV attention, respectively, and assess their impacts on model performance by comparing with the baseline that has the same number of KV heads. Based on the results shown in Table 1, in most cases, reducing the number of K heads has a relatively minor impact on model performance compared to reducing the number of V heads. This differential impact is reasonable, considering the distinct roles played by the K and V components within the attention mechanisms. The K heads primarily serve to compute attention matrices, which prove to have a remarkable sparsity exceeding 95% during inference (Zhang et al., 2024b). This high sparsity nature implies that a slight decrease in the parameters used for its calculation can still yield a precise approximation. In contrast, the V vectors, which directly influence the final attention output, demands a more nuanced approach.
@@ 2.2 Observations
Observation 2
Reducing the dimension of K heads to half of the dimension of V heads ($d^{h}_{k}=d^{h}_{v}/2$) can preserve the model performance, compared to the scenario where the dimensions of K and V heads are equal ($d^{h}_{k}=d^{h}_{v}$). In Observation 1, we demonstrate the minimal effects resulting from a decrease in the number of K heads. Here, we explore an alternative approach to compressing K components by reducing the dimension of K heads. More specifically, “the dimension of K heads” refers to $d^{h}_{k}$ in § 2 and its reduction indicates a smaller $W_{K}$ and $k_{t}$). To compute the attention scores when the dimensions of K and V are different (Eq. 8), we transform the dimension of $K$ to the same as $V$ through a trainable feed-forward layer. In essence, the reduction of K head dimension reduces KV cache at the cost of additional computation from a feed-forward layer. This trade-off can be very cost-effective as the primary bottleneck for LLM inference speed lies in memory consumption rather than computation (Shazeer, 2019; Ribar et al., 2024). To investigate the impact of reducing K head dimension, we conduct a comparison of various attention settings, such as MHA and GQA, along with their counterparts having only half of the K head dimension. The results presented in Table 2 demonstrate that the decrease in performance due to reducing the K head dimension is negligible. In certain configurations, the model with a smaller K head dimension even outperforms the baseline setting prior to compression.
Model Overall Commonsense & Comprehension Continued LM
Hella. ObQA Wino. ARC. PIQA SciQ Bool. Logi. LAMB.
GQA ($n_{k}^{h}$=$n_{v}^{h}$=4) 51.66 54.0 38.0 56.0 37.5 72.3 82.0 61.3 28.6 35.4
Observation 3
Using only a small proportion of V vectors, specifically those corresponding to the highest attention scores, in the attention output calculation can still preserve the model performance. As in previous work, topk V can improve the memory efficiency of the model (Gupta et al., 2021), inspiring us to adopt this strategy in this KV imbalance scenario. As shown in Table 3, even when we only use 100 V vectors corresponding to the top100 highest attention scores (denoted as “Sel.V-top100”), the model can still maintain a comparable effect to the original setting. These results indicate that we only need to retrieve a small proportion of V vectors from the cache during inference to achieve competitive performance. It significantly reduces data transfer, and thereby improves inference speed.
@@
Observation 4
Augmenting Q components can boost model performance. Augmenting the self-attention layer with extra parameters is a straightforward way to enhance its representational capacity and improve model performance. Since the Q vectors in self-attention do not need to be cached during inference, their impact on memory usage and data transfer duration is minimal. Therefore, investing extra parameters in the Q components can be more cost-effective compared to the KV components. In the following, we investigate whether augmenting Q components can boost model performance through experiments. In our experiments, we varied the number of parameters in Q to assess their impact on model performance. As demonstrated in Table 4, adding extra parameters to Q can enhance the model’s performance to different extents, especially when the compression levels of K and V are higher ($n^{h}_{v}$=4,$n^{h}_{k}$=4). Based on the overall model performance, a scaling factor of 1.5$\times$ (i.e. $d_{q}^{h}=3072$) seems to be optimal.
@@
2.3 Summary: Optimal Configuration for DiffQKV Attention
Based on the above observations, we can summarize the optimal configurations for DiffQKV attention that can well balance model performance and inference efficiency as follows: 1) differentially compressed KV: unlike compressing V components, compressing K incurs less performance degradation, whether in terms of the number of K heads ($n^{h}_{k}$) or the dimension of K ($d^{h}_{k}$); Furthermore, for V with less compression, we can reduce its memory cost by employing the selective V cache fetching. 2) augmented Q: introducing extra parameters to the Q components can enhance the representational capacity of the self-attention layer, while having a relatively small impact on inference efficiency, since the Q vectors do not need to be cached during inference. In Table 6, by integrating three aforementioned strategies (i.e., -75% K Heads ($n_{k}^{h}$=4), Half K Dim($d^{h}_{k}$= $d^{h}_{v}/2$), and AugQ ($d_{q}^{h}$=3072), we observe that the model’s performance surpasses that of the original model and also exhibits advantages in terms of inference speed and KV cache utilization.
In summary, by employing an imbalanced KV head, we have developed a more efficient model architecture. Furthermore, the introduction of the augmented Q module allows us to compensate for the loss incurred by compression while achieving superior performance compared to standard models. The implications of this approach extend beyond the Sigma model, particularly in light of recent trends toward increasing model capacity. To reduce the KV cache cost during inference in large models, there has been a gradual trend toward reducing the proportion of attention mechanism parameters within the entire Transformer block. Given the critical role of the attention mechanism in language models, its compression cannot be done without limits. By reinforcing the Q head that does not require caching, our method introduces new possibilities for scaling up the attention mechanism, thereby enriching its representational space. Table 5 highlights the parameter efficiency gained from scaling up the parameters of the attention mechanism. However, incorporating the augmented Q inevitably introduces additional inference costs. We will further elucidate the practical efficiency of Sigma in § 3.
2.4 Sigma Model Architecture
Building on the insights of the DiffQKV attention discussed in § 2.3, we construct Sigma, a pre-trained language model characterized by strong performance and inference efficiency. Specifically, we adopt two model scales with 1.5 billion parameters and 10 billion parameters, respectively (i.e. Sigma-1.5B and Sigma-10B). For the sake of balancing the model performance and the cost of the KV cache, during the training of Sigma-1.5B and Sigma-10B, no dimension compression is applied to the K heads. Only the number of K heads was decreased. Concretely, the K head is set to 4, and the number of V heads, which is half of the number of Q heads, is set to 16. For Sigma-1.5B, we set $d_{q}^{h}=3072$, and for Sigma-10B, we set $d_{q}^{h}=6144$, corresponding to 1.5 times the dimension of the hidden state, so as to extend the representational space of Q. We utilize the same vocabulary as Llama3 (Dubey et al., 2024), with a vocabulary size of 128k. For more detailed configurations of the Sigma architecture, please refer to Table 7.
Table 7: Key configurations and hyperparameters of Sigma 1.5B and 10B.
@@ 2.4 Sigma Model Architecture
3 Efficiency Analysis
In this section, we conduct a detailed efficiency analysis of Sigma-1.5B, using flash attention 2 (Dao et al., 2022; Dao, 2024) as the default attention implementation. We provide both theoretical and empirical evaluations to explore the efficiency gains achieved by Sigma. Through a detailed examination of its design and performance across multiple metrics, we aim to highlight the architectural innovations and optimization techniques that underpin its superior efficiency in addressing complex and diverse tasks.
3.1 Theoretical Analysis
As outlined in § 2.3, the efficiency gains of Sigma can be largely attributed to the reduction in the number of key heads. In Sigma 1.5B, the number of key heads is reduced from 16 to 4, while maintaining the same number of value heads. This reduction directly impacts two critical components of the attention layer: KV Cache and Attention Computation. Below, we delve into the implications and benefits of this architectural adjustment.
KV Cache.
[…]
$L=\alpha\cdot[\texttt{bsz}\cdot\texttt{seq\_len}\cdot(n^{h}_{k}\cdot d^{h}_{k}+n^{h}_{v}\cdot d^{h}_{v})]+\beta.$ (14)
[…]
$r=\lim_{\texttt{seq\_len}\rightarrow+\infty}\frac{L^{\rm GQA}-L^{\rm{Sigma}}}{L^{\rm GQA}}=\frac{32-20}{32}=37.5\%.$ (15)
Attention Computation.
Here, the cost reduction also originates from the reduction in key heads. The theoretical reduction rate $r$ therefore remains 37.5%. However, while KV cache operations are primarily I/O-intensive, attention computation is significantly more computation-intensive. Due to the increased complexity of this module, the observed reduction in practice is likely to deviate more significantly from the theoretical rate compared to the KV cache.
Above, we discussed the theoretical efficiency improvements of Sigma compared to the vanilla GQA design. Additionally, it is important to highlight that the reduction in the size of the key cache also decreases the overall memory consumption for the KV cache. As a result, Sigma allows for larger batch sizes, which can naturally enhance the efficiency of inference.
3.2 Implementation
Although the reduction in the number of key heads theoretically leads to significant efficiency improvements, we find it challenging to implement in practice due to the high integration of KV cache and attention operations within flash attention and existing LLM deployment frameworks (e.g., vLLM (Kwon et al., 2023) and TensorRT [^2] https://github.com/NVIDIA/TensorRT-LLM/). In this section, we propose several temporary solutions for deploying Sigma and highlight the need for broader support for DiffQKV.
Load and Storage of KV Cache.
In Sigma, since key and value have different number of heads, their cache sizes differ. Generally, it poses no issues for most implementations that store KV cache separately. However, for frameworks like vLLM, it becomes challenging, as these kinds of frameworks combine key cache and value cache into a whole matrix. To accommodate these frameworks, one workaround is to duplicate the key matrix to match the size of the value matrix (named KV Group Sharing). However, this approach is highly discouraged, as it negates the efficiency improvements entirely while merely maintaining the same performance as the original design. We release the official implementation of KV Group Sharing in Appendix E.
Flexible Attention Computation.
The official Flash Attention only supports attention computations when the number of key heads equals the number of value heads, and the number of query heads is an integer multiple of the number of key and value heads. Therefore, we could not directly use Flash Attention in Sigma. To address this issue, we introduce a flexible version of FlashAttention, named FlexHeadFA [^3] https://github.com/xiayuqing0622/flex_head_fa, which leverages address probing to support attention computation with a flexible number of heads. Specifically, FlexHeadFA decouples the address calculations for key heads and value heads. For each query head index $\texttt{idx}_{q}$, it calculates the corresponding key head index $\texttt{idx}_{k}$ and value head index $\texttt{idx}_{v}$ using equations 16 and 17. This approach enables FlexHeadFA to retrieve the appropriate elements during element-wise multiplication using the calculated indices, thereby eliminating the constraint of requiring equal head numbers.
$\texttt{idx}_{k}=\texttt{idx}_{q}\cdot\frac{n^{h}_{k}}{n^{h}_{q}},$ (16)
$\texttt{idx}_{v}=\texttt{idx}_{q}\cdot\frac{n^{h}_{v}}{n^{h}_{q}}.$ (17)
More precisely, FlexHeadFA’s implementation follows that of Flash Attention 2, consisting of flash_fwd_splitkv_kernel and flash_fwd_splitkv_combine_kernel. These are two major GPU kernel operations. The former, referred to as the split kernel, manages the primary attention computation workload by dividing the key and value matrices into smaller chunks and performing multiplications to generate output chunks. Meanwhile, the latter, referred to as the combine kernel, assembles these output chunks into a complete output matrix.
3.3 Empirical Analysis
In this section, we conduct comprehensive experiments to validate and substantiate our theoretical estimations of efficiency improvements over the vanilla GQA design.
3.3.1 Experiment Setup
Model Settings.
To demonstrate the efficiency gains achieved by reducing the number of key heads, we compare Sigma 1.5B with a hypothetical baseline model, referred to as the standard model (Std). To be specific, while Sigma features unbalanced KV heads ($n_{k}^{h}=4$ and $n_{v}^{h}=16$) and includes an additional augmented Q ($d_{q}^{h}=3072$), Std employs a standard Group Query Attention (GQA) mechanism, configured with balanced K and V heads ($n_{k}^{h}=16$ and $n_{v}^{h}=16$).
Metric Settings.
Two methods are employed to accurately measure the efficiency improvements: Cuda Event Elapsed Time and Kernel Execution Time. We also record the absolute and relative improvement of these metrics.
  • Cuda Event Elapsed Time places checkpoints (torch.cuda.Event) within the code (modeling_sigma.py) and measures the elapsed time between them to estimate the total duration of specific operations. Its primary advantage lies in its high flexibility, allowing developers to record any time interval by modifying the code. However, it would be unstable when measuring operations with small absolute time scales. We offer an official implementation in Appendix F for reference.
  • Kernel Execution Time provides a more precise time estimate by profiling GPU applications directly and recording the execution times of GPU kernels using nsys. While this method ensures high accuracy, it incurs significant overhead as the inference scale increases. Additionally, capturing certain kernel operations requires substantial expertise.
Task Settings.
Since the above two metrics have separate application scopes, we assign different tasks to them, conducting all experiments on a NVIDIA H100 80G HBM3 GPU.
  • Kernel Execution Time (KET Test): In this experiment, we measure the cost of two kernels: flash_fwd_splitkv_kernel and flash_fwd_splitkv_combine_kernel, treating their summation as the overall cost of applying FlexHeadFA. As this metric is designed for small-scale recording, both models (Sigma 1.5B and Std) are tasked with generating 10 tokens while varying the prefix length in a grid pattern: [2k, 4k, 16k, 32k]. Due to CUDA memory limitations caused by nsys, we were unable to record KET for a 64k prefix. We compare the cumulative value of these metrics.
  • Cuda Event Elapsed Time (CEET Test): In this experiment, we record the cost of three components: the cost of storing and loading KV Cache, applying Attention Computation and applying Augmented Q. Although Std does not principally include the Augmented Q operation, we incorporate it additionally and compare its performance to ensure that the efficiency comparison is not influenced by variations in machine status. Both models are evaluated on larger-scale inference tasks, with the output length increasing in a grid pattern of [2k, 4k, 8k, 16k, 32k, 64k], and the prefix length varying independently as [0, 2k, 4k, 16k, 32k, 64k]. We compare the cumulative value of these metrics.
3.3.2 Empirical Results
(a) KET of the split kernel.
(b) KET of the combine kernel.
(c) Total KET.
@@ 3.3.2 Empirical Results
KET Results.
We present Kernel Execution Time results in Figures 2 and 8, which align closely with our previous analyses. The cost flash_fwd_splitkv_kernel constitutes the majority of the computational time and increases sharply as the prefix length grows. In contrast, the cost of flash_fwd_splitkv_combine_kernel rises at a much slower rate, gradually accounting for a smaller proportion of the overall computation time. Regarding relative improvement, the efficiency gains become increasingly pronounced as the prefix length increases. We anticipate that the absolute improvement value will gradually approach 37.5% as the prefix length continues to grow. Due to GPU memory constraints, we were unable to further validate this assertion using this metric.
Examining the individual improvement of the separate kernels, flash_fwd_splitkv_kernel, which directly operates on the key and value matrices, benefits naturally from the reduction in key heads. This reduction decreases the memory load for the key matrix, and as a result, the improvement ratio for this kernel is expected to converge to 37.5% with increasing prefix length. In contrast, the combination of output chunks is less associated with the reduction of key heads. Hence, flash_fwd_splitkv_combine_kernel, responsible for combining output chunks, is less influenced by the reduction in key heads. Consequently, its improvement ratio remains consistently near zero, indicating that this operation is unaffected by changes in the key matrix. However, as we report cumulative values, the absolute value naturally increases alongside the prefix length.
Here, we report the Cuda Event Elapsed Time for three modules: KV Cache, Attention Computation, and Augmented Q, with the results presented in Figures 5, 4, 6 and 3.
@@
CEET Results - Augmented Q.
@@
CEET results of KV cache can be found in Figure 4. CEET here primarily reflects the cost of loading and storing the KV cache. These operations are largely proportional to the size of the key and value matrices with minimal extraneous influencing factors. Therefore, the results for KV cache align most closely with our theoretical analysis. In Figures 4(a), 4(b), 4(c), 4(d), 4(e) and 4(f), Sigma demonstrates a significantly lower cost for KV cache operations. As the output length increases from Figure 4(a) to Figure 4(e), the CEET gap between SIGMA and Std in KV cache becomes increasingly pronounced. As illustrated in Figure 4(h), when the prefix and output lengths reach 64k, Sigma achieves a speedup of 36.57% compared to Std, which is very close to the theoretically calculated improvement rate of 37.5%.

### v2

@@ Appendix B Detailed Illustration of DiffQKV Attention
Model Overall Commonsense & Comprehension Continued LM
Hella. ObQA Wino. ARC. PIQA SciQ Bool. Logi. LAMB.
GQA ($n_{k}^{h}$=$n_{v}^{h}$=4) 51.66 54.0 38.0 56.0 37.5 72.3 82.0 61.3 28.6 35.4
B.2 Selective V Cache Fetching
Apart from differentially compresssed KV and augmented Q, we can further optimize this part by loading V vectors selectively during inference. Due to the high sparsity exhibited by the attention scores (Xiao et al., 2023; Zhang et al., 2024b), it can still preserve the model performance with only a small number of V vectors to approximate the results, while largely reducing the memory usage. We term this strategy as selective V cache fetching (Xiao et al., 2023; Zhang et al., 2024b). Due to the high sparsity in attention scores, efficiency can be further enhanced by loading V vectors selectively during inference.
@@ B.2 Selective V Cache Fetching
Specifically, we observe that using only a small proportion of V vectors, specifically those corresponding to the highest attention scores, in the attention output calculation can still preserve the model performance. As in previous work, topk V can improve the memory efficiency of the model (Gupta et al., 2021), inspiring us to adopt this strategy in this KV imbalance scenario. As shown in Table 8, even when we only use 100 V vectors corresponding to the top100 highest attention scores (denoted as “Sel.V-top100”), the model can still maintain a comparable effect to the original setting. These results indicate that we only need to retrieve a small proportion of V vectors from the cache during inference to achieve competitive performance. It significantly reduces data transfer, and thereby improves inference speed.
@@
B.3 Detailed Evaluation Results
Due to the space limitation of the main text, in Table 3-4, we only include the average results across nine benchmarks. Here, we report the detailed evaluation results on each benchmark in Tables 9, 10, 11 and 12.
B.4 Combination of Multiple Optimization Strategies
Table 14: Key configurations and hyperparameters of Sigma.
@@ B.4 Combination of Multiple Optimization Strategies
As shown in Table 13, by integrating three optimization strategies (i.e., -75% K Heads ($n_{k}^{h}$=4), Half K Dim($d^{h}_{k}$= $d^{h}_{v}/2$), and AugQ ($d_{q}^{h}$=3072), we observe that the model’s performance surpasses that of the original model and also exhibits advantages in terms of inference speed and KV cache utilization.
B.5 Sigma Model Architecture
Building on the DiffQKV attention, we construct a pre-trained language model, named Sigma. Specifically, we adopt two model scales with 1.5B parameters and 10B parameters, respectively (i.e. Sigma-1.5B and Sigma-10B). For the sake of balancing the model performance and the cost of the KV cache, during the training of Sigma-1.5B and Sigma-10B, no dimension compression is applied to the K heads. Only the number of K heads was decreased. Concretely, the K head is set to 4, and the number of V heads, which is half of the number of Q heads, is set to 16. For Sigma-1.5B, we set $d_{q}^{h}=3072$, and for Sigma-10B, we set $d_{q}^{h}=6144$, corresponding to 1.5 times the dimension of the hidden state, so as to extend the representational space of Q. We utilize the same vocabulary as Llama3 (Dubey et al., 2024), with a vocabulary size of 128k. For more detailed configurations of the Sigma architecture, please refer to Table 14.
Appendix C Evaluation Procedure on AIMicius Benchmark
C.1 CMDGen
We consider the following evaluation metrics on the CMDGen task:
  • CMD Score computes the cosine similarity between the embeddings of the generated command and the ground-truth command. The embeddings are encoded with all-MiniLM-L6-v2.555https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2
C.2 Infrawise
We adopt the following metrics to conduct evaluation on Infrawise:
C.3 Optiflow
C.4 NL2KQL
    cluster(CLUSTER).database(DATABASE).TABLE
    | where COLUMN
Appendix D Data Examples of AIMicious Benchmark
D.1 CMDGen Examples
D.2 NL2KQL Examples
D.3 Infrawise Examples
D.4 Optiflow Examples
Appendix E Official Implementation of KV Group Sharing
1 """
2␣␣␣␣ThisisforbalancethenumberofKVheadstofacilitatesubsequentflashattentioncalculations
3␣␣␣␣"""
4 batch, _, slen, head_dim = key_states.shape
5
6 if self.num_key_heads > self.num_value_heads:
7 n_rep = int(self.num_key_heads / self.num_value_heads)
10 else:
11 n_rep = int(self.num_value_heads / self.num_key_heads)
12 if n_rep == 1:
13 return key_states, value_states
You can use this code in the forward function like this:
1 self,
2 hidden_states: torch.Tensor,
3 attention_mask: Optional[torch.LongTensor] = None,
4 position_ids: Optional[torch.LongTensor] = None,
5 past_key_value: Optional[Cache] = None,
6 output_attentions: bool = False,
7 use_cache: bool = False,
8 cache_position: Optional[torch.LongTensor] = None,
9) -> Tuple[torch.Tensor, Optional[torch.Tensor], Optional[Tuple[torch.Tensor]]]:
10 if isinstance(past_key_value, StaticCache):
11 raise ValueError(
14 )
15 output_attentions = False
16 bsz, q_len, _ = hidden_states.size()
17
18 query_states = self.q_proj(hidden_states)
19
20 key_states = self.k_proj(hidden_states)
21
23
24 value_states = self.v_proj(hidden_states)
25
26 # Flash attention requires the input to have the shape
27 # batch_size x seq_length x head_dim x hidden_dim
28 # therefore we just need to keep the original shape
29
33
34 cos, sin = self.rotary_emb(value_states, position_ids)
36
37 if past_key_value is not None:
41
43 query_states = query_states.transpose(1, 2)
44 key_states = key_states.transpose(1, 2)
45 value_states = value_states.transpose(1, 2)
46
47 (Code Omitted.)
48
49 ’’
50␣␣␣␣flashattention
51␣␣␣␣’’
52 attn_output = self._flash_attention_forward(
54 )
55
56 attn_output = attn_output.reshape(bsz, q_len, -1).contiguous()
57 attn_output = self.o_proj(attn_output)
58
59
60 if not output_attentions:
61 attn_weights = None
62
63 return attn_output, attn_weights, past_key_value
Appendix F Official Implementation of Efficiency Recording
Here, we offers an official implementation to use Cuda Event Elapsed Time to record the cost of attention computation in modeling_sigma.py.
1start = torch.cuda.Event(enable_timing=True)
2end = torch.cuda.Event(enable_timing=True)
3start.record()
4attn_output = flash_attn_func(
6)
7end.record()
8end.synchronize()
9latency = start.elapsed_time(end)
10total_attention_cost += latency
Appendix G Detailed Efficiency Results
@@
Here, we report detailed KET results and CEET results of three modules: KV Cache, Attention Computation, and Augmented Q, as shown in Figures 6, 5, 4 and 15.
@@
CEET results of KV cache can be found in Figure 5. CEET here primarily reflects the cost of loading and storing the KV cache. These operations are largely proportional to the size of the key and value matrices with minimal extraneous influencing factors. Therefore, the results for KV cache align most closely with our theoretical analysis. In Figures 5(a), 5(b), 5(c), 5(d), 5(e) and 5(f), Sigma demonstrates a significantly lower cost for KV cache operations. As the output length increases from Figure 5(a) to Figure 5(e), the CEET gap between SIGMA and Std in KV cache becomes increasingly pronounced. As illustrated in Figure 5(h), when the prefix and output lengths reach 64k, Sigma achieves a speedup of 36.57% compared to Std, which is very close to the theoretically calculated improvement rate of 37.5%.